OpenStudy (gamereternity):

What is the ratio 60/135 written in word form and in lowest terms? A. 6 to 17 B. 15 to 32 C. 4 to 9 D. 10 to 13

Directrix (directrix):

Can you think of any factors common to both 60 and 135 ?

OpenStudy (gamereternity):

Sorry forgot to put im checking my answers im pretty sure c

OpenStudy (gamereternity):

@Directrix

Directrix (directrix):

Let's see. 60 = 15*4 135 = 15*9

Directrix (directrix):

What about this part of the question: >> written in word form Are you supposed to give 4/9 in word form?

Directrix (directrix):

@gamereternity

OpenStudy (gamereternity):

no

Directrix (directrix):

I looked again at the options and 4 to 9 is the "word" form of 4/9. If you need a ratio study site, this is fairly good: https://www.mathsisfun.com/numbers/ratio.html